Solve the equation for r. h= Vr^-2/3.14 I think I might have this figured out, but I'm not certain it is correct. Here is the final answer I got. r=(1/pie symbol*h)/V. All raised to the power of 2
Answer by scott8148(6628)

h = V / (r^2 * pi)

r^2 = V / (h * pi)

r = sqrt[V / (h * pi)]